Solve for x and y in the following set of simultaneous differential equations by using D-operator methods: (D-2)x + Dy = 10sin2t

Dx + (D+2)y = 0


1
Expert's answer
2021-08-10T12:56:18-0400

We know that the operator "D" means a derivative "D=\frac{d}{dt}" , but for now we will assume that this is just another letter and solve the specified system of equations by the "addition" method.



"\left\{\begin{array}{l}
\left.\left(D-2\right)x+Dy=10\sin2t\right|\times\left(D\right)\\[0.3cm]
\left.Dx+\left(D+2\right)y=0\right|\times\left(D-2\right)
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
D\left(D-2\right)x+D^2y=D\left(10\sin2t\right)\\[0.3cm]
\text{Subtract the lower one from the upper equation}\\[0.3cm]
D\left(D-2\right)x+\left(D^2-4\right)y=0
\end{array}\right.\\[0.3cm]\\[0.3cm]
\left(\cancel{D^2}-\cancel{D^2}+4\right)y=D\left(10\sin2t\right)\\[0.3cm]
4y=\frac{d}{dt}\left(10\sin2t\right)=20\cos2t\to\boxed{y(t)=5\cos2t}\\[0.3cm]
\text{To find the function} \quad x(t), \text{go back to the original system.}\\[0.3cm]
\left\{\begin{array}{l}
\left.\left(D-2\right)x+Dy=10\sin2t\right|\times\left(D+2\right)\\[0.3cm]
\left.Dx+\left(D+2\right)y=0\right|\times\left(D\right)
\end{array}\right.\\[0.3cm]
\left\{\begin{array}{l}
\left(D^2-4\right)x+D\left(D+2\right)y=\left(D+2\right)\left(10\sin2t\right)\\[0.3cm]
\text{Subtract the lower one from the upper equation}\\[0.3cm]
D^2x+D\left(D+2\right)y=0
\end{array}\right.\\[0.3cm]\\[0.3cm]
\left(\cancel{D^2}-\cancel{D^2}-4\right)x=\left(D+2\right)\left(10\sin2t\right)\\[0.3cm]
-4x=\left(\frac{d}{dt}+2\right)\left(10\sin2y\right)=20\cos2t+20\sin2t\to\\[0.3cm]
\boxed{x(t)=-5\cos2t-5\sin2t}\\[0.3cm]"

ANSWER



"\left\{\begin{array}{l}
x(t)=-5\cos2t-5\sin2t\\[0.3cm]
y(y)=5\cos2t
\end{array}\right."